### v2

@@ 5 Experimental comparison of ranking functions
In this section, we present some proportional ranking functions with concave activation functions and compare them using no-regret dynamics simulations. We highlight that our goal is to study the long-term behavior and dynamics of strategic, rational agents that engage with algorithmic tools for regret minimization, which is different than (and complementary to) other lines of work that evaluate the dynamics of human publishers that may be prone to various biases, e.g. [RRTK17]. Therefore, we chose to evaluate the ranking functions using a code simulation in which agents use the state-of-the-art no-regret algorithm of [FAL+22], which is designated for concave games as in our case. We compare three families of proportional ranking functions based on some concave and decreasing functions as activation functions.
  • The linear proportional ranking function family: $g_{lin}(t)=b-t$, where $b>1$ is the intercept hyperparameter.
  • The root proportional ranking function family: $g_{root}(t)=(1-t)^{a}$, where $10<a<$ 1 is the power hyperparameter.
  • The logarithmic proportional ranking function family: $g_{log}(t)=\ln(c-t)$, where $c>2$ is the shift hyperparameter.
Let $r^{lin}$, $r^{root}$ and $r^{log}$ be the proportional ranking functions induced by these activation functions. Unless stated otherwise, we use the default values of $b=1+\delta$, $a=\frac{1}{2}$, $c=2+\delta$, where $\delta\coloneqq 10^{-5}$.
5.1 Simulation details
Our simulations utilize the state-of-the-art no-regret algorithm presented in [FAL+22], the Log-Regularized Lifted Optimistic FTRL (hereafter, LRL-OFTRL). For each player, the algorithm uses the current gradient of the utility, the gradients from previous rounds, and a learning rate $\eta$ to decide which action to play in the next round. The important feature of LRL-OFTRL is the logarithmic increase in regret as the dynamics progresses. In our simulations, which are carried out in rounds, we define a publishers’ game $G$ and let the players choose their action in each round using the LRL-OFTRL algorithm. A simulation is said to converge if it has reached an $\varepsilon$-NE. Throughout all simulations, we used homogeneous $\lambda_{i}=\lambda$ for all $i\in N$, the squared Euclidean semi-metric $d(x,y)=\frac{1}{k}||x-y||_{2}^{2}$, learning rate $\eta=1/2$ and $\varepsilon=10^{-4}$. In all simulations, we consider uniform demand distributions over a sample of $s\in\operatorname*{\mathbb{N}}$ information needs.666 Note that while we restrict our simulations to uniform demand distributions, any demand distribution can be approximated with $s$ large enough.
@@ 5.1 Simulation details
To evaluate the performance of the different ranking functions, we report three evaluation criteria: the publishers’ welfare, the users’ welfare and the convergence rate. The two welfare measures are defined similarly to previous work [MPRT24, YLS+24]. The publishers’ welfare is defined to be the sum of publisher utilities:
$\mathcal{U}(x)\coloneqq\sum_{i=1}^{n}u_{i}(x)=1-\lambda\sum_{i=1}^{n}d^{0}_{i}(x_{i})$ (4)
@@
These welfare measures can be evaluated at any strategy profile $x\in X$, but we report the values obtained in the $\varepsilon$-NE the dynamics converged to, as these values represent the long-term welfares. The convergence rate is simply the number of rounds it took the dynamics to converge. For each parameter configuration and each evaluation criterion, we performed $500$ simulations, when initial documents and demand distributions were drawn uniformly i.i.d., and constructed bootstrap confidence intervals with a confidence level of 95% using a bootstrap sample size of $B=500$.
5.2 Simulations results
5.2.1 Comparison between ranking function families
Figure 1 presents several interesting trends. A prominent phenomenon in the figure is a trade-off between the different measures - whenever an activation function induces an equilibrium in which the publishers’ welfare is high, the users’ welfare is lower and the game dynamics converge more slowly. The activation functions that yield higher publishers’ welfare values are the ones that align the ranking function more closely with a uniform ranking (i.e., $r_{i}\equiv\frac{1}{n}$), or, in other words, the ones in which the impact of the publishers’ content on the ranking is less significant. When ranking the documents using these functions, the publishers are less incentivized to create relevant content and thus stay closer to their respective initial documents. This leads to higher publishers’ welfare and lower users’ welfare. On the contrary, there are less uniform ranking functions, in which the distances from the information needs have a greater influence on the ranking. Those ranking functions encourage publishers to compete for each information need and to generally make their documents more relevant to the users, which leads to higher users’ welfare and lower publishers’ welfare. [^7] To elucidate this point, let us compare the activation functions $g_{lin}$ and $g_{root}$. Note that $g_{root}$ is a concave transformation of the linear activation function. When applying a square root to a number in $[0,1]$, the smaller it is the bigger the factor it is increased by. Therefore, for any fixed profile, the ranking produced by $g_{root}$ is closer to uniform distribution than that produced by $g_{lin}$. In addition, when the ranking functions are less uniform, the absolute values of the gradients increase, which leads to faster convergence.
In the rest of the section, we demonstrate that this order relation between the ranking functions is robust to changes in the ecosystem structure, such as changes in the game parameters, and show how these changes affect our evaluation criteria.
The effect of the penalty factor $\lambda$
Figure 1 illustrates how $\lambda$, the penalty factor, influences our evaluation criteria, showing similar trends to those discussed in [MPRT24]. One might expect an increase in $\lambda$ to harm the publishers’ welfare, since a larger $\lambda$ imposes greater penalty on the publishers for providing content that differs from their initial documents. Yet, a higher $\lambda$ also incentivizes the publishers to adhere more closely to their initial documents, which increases the publishers’ welfare. This two opposing forces explain the non-monotonic trend observed in Figure 1 - for lower values of $\lambda$, the publishers’ welfare decreases as $\lambda$ increases, until it reaches a minimum point, after which it keeps increasing. The trend in the users’ welfare is simpler: it monotonically decreases as $\lambda$ increases, likely because higher penalties lead to less relevant content from publishers. Lastly, there is no significant impact of the penalty factor on the convergence rate.
@@ 5.2.1 Comparison between ranking function families
As depicted in Figure 3, increasing $s$ leads to an increase in the publishers’ welfare and a decrease in the users’ welfare, both trends eventually plateauing. A possible explanation is that when increasing $s$ while fixing $n$, a fixed number of publishers compete for an increasing number of information needs, which makes it harder for the publishers to maximize their exposure. Recall that a publisher’s utility from exposure is the expected value of her exposure for different information needs, so when a player is unable to meet a large portion of the information needs, she might be demotivated to try to be relevant to the users and focus on being faithful to her initial document, boosting the publishers’ welfare in the expense of the users.
The convergence rate is affected by $s$ as well. When $s$ is small, the publishers must compete for the same information needs, which leads to more complicated dynamics and hence to slower convergence. As $s$ increases, the publishers tend to focus on different information needs (the ones which are the closest to their initial document), therefore we get much simpler dynamics and faster convergence.
@@
5.2.2 Comparison within ranking function families
Beyond the comparison between the three ranking function families, tuning the hyperparameters within each family also reveals interesting phenomena. The experimental results supporting these observations are provided in Appendix B, and are discussed here. The results once again underscore the publisher-user trade-off discussed in §5.2.1: hyperparameter values that lead to higher publishers’ welfare result in lower users’ welfare and slower convergence rate, and vice versa. As was the case in the comparison between ranking function families, the hyperparameters’ values that lead to higher publishers’ welfare (and lower users’ welfare) are the values that make the ranking function closer to uniform ranking. For example, a decrease in the power parameter $a$ of $r^{root}$ can be seen in Figure 6 to correlate with an increase in publishers’ and a decrease in users’ welfare, and indeed when the power parameter $a$ approaches 00, the root proportional ranking function approaches a uniform ranking. That is: $r^{root}_{i}\to\frac{1}{n}$ as $a\to 0$. Similarly, both the intercept parameter $b$ of $r^{lin}$ and the shift parameter $c$ of $r^{log}$ are positively correlated with the publishers’ welfare and negatively correlated with the users’ welfare, and indeed $r^{lin}_{i}\to\frac{1}{n}$ as $b\to\infty$ and $r^{log}_{i}\to\frac{1}{n}$ as $c\to\infty$.
@@ 6 Discussion
We studied an information retrieval game in which publishers have initial documents they prefer to provide and initiated the analysis of no-regret dynamics in such games. We defined the class of proportional ranking functions, which are ranking functions that are determined by a one-variable scalar activation function, and established a full characterization of which proportional ranking functions induce socially-concave games and which induce concave games. We then used this characterization to conclude that any concave activation function guarantees the convergence of no-regret dynamics, thereby offering a system designer a rich family of ranking functions that, under the assumption that publishers minimize their regret, guarantee the reach of the system to a stable state. We then empirically investigated the publishers’ welfare and the users’ welfare, as well as the convergence rate, which measures how long it takes the ecosystem to reach stability. In particular, we examined how these three criteria are affected by the choice of the ranking function and by changes in the ecosystem structure.
Limitations
Although we prove convergence of no-regret dynamics whenever the activation function is concave, our approach provides no insights regarding convergence in the case it is not concave. Further, we assumed each publisher is restricted to publishing only one document, while in practice a publisher can write several documents, e.g. in order to compete for exposure from multiple information needs simultaneously. A specific cost function was assumed to capture the cost the publishers incur for deviating from their initial documents. A major limitation of our experiments is that we drew $x_{0}$ and $x^{*}$ i.i.d from a uniform distribution, and the trends presented in our analysis are not guaranteed to hold for other ecosystem distributions. A theoretical welfare and convergence rate analysis are also out of scope. We believe these limitations can serve as interesting directions for future work, and further expand our understanding of the dynamics of strategic publishers.
@@ Acknowledgments
This work was supported by funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement 740435).
Appendix A Proofs
A.1 Proof of Lemma 1
@@
since $\sum_{i=1}^{n}r_{i}(x;x^{*})=1$ for any strategy profile $x\in X$ and information need $x^{*}\in[0,1]^{k}$. Now, $g(x)$ is concave since for every $i\in N$ and $x_{0}^{i}\in[0,1]^{k}$, $d(x_{i},x_{0}^{i})$ is convex in $x_{i}$. Regarding condition A2., it can be easily shown (using the linearity and monotonicity of expectation) that the convexity of $r_{i}(x;x^{*})$ in $x_{-i}$ for any $x_{i}\in X_{i},x^{*}\in[0,1]^{k}$ implies that $\mathbb{E}_{x^{*}\sim P^{*}}\left[r_{i}(x;x^{*})\right]$ is convex in $x_{-i}$ for any $x_{i}\in X_{i}$. Together with the fact that $d(x_{i},x_{0}^{i})$ is fixed in $x_{-i}$, we get that $u_{i}(x)=\mathbb{E}_{x^{*}\sim P^{*}}\left[r_{i}(x;x^{*})-\lambda_{i}\cdot d(x_{i},x_{0}^{i})\right]$ is convex in $x_{-i}$ for any $x_{i}\in X_{i}$. ∎
A.2 Full derivation of inequality $f^{\prime\prime}(\hat{a})>0$ from the proof of Theorem 1
This appendix provides a detailed derivation of the equation $f^{\prime\prime}(\hat{a})>0$ from the proof of the main result. By the definition of the utility in publishers’ games, and since in the game we constructed the distribution over the information needs is a one-hot distribution that places probability $1$ on $x^{*}=0$:

### v3

@@ 5 Experimental comparison of ranking functions
In this section, we present some PRFs with concave activation functions and compare them using no-regret dynamics simulations. We highlight that our goal is to study the long-term behavior and dynamics of strategic, rational agents that engage with algorithmic tools for regret minimization, which is different from (and complementary to) other lines of work that evaluate the dynamics of human publishers that may be prone to various biases, for example, Raifer et al. (2017). Therefore, we chose to evaluate the ranking functions using a code simulation in which agents use the state-of-the-art no-regret algorithm of Farina et al. (2022), which is designated for concave games as in our case. We compare three families of PRFs based on some concave and decreasing functions as activation functions.
  • The linear PRF family: $g_{lin}(t)=b-t$, where $b>1$ is the intercept hyperparameter.
  • The root PRF family: $g_{root}(t)=(1-t)^{a}$, where $10<a<$ 1 is the power hyperparameter.
  • The logarithmic PRF family: $g_{log}(t)=\ln(c-t)$, where $c>2$ is the shift hyperparameter.
Let $r^{lin}$, $r^{root}$ and $r^{log}$ be the PRFs induced by these activation functions. Unless stated otherwise, we use the default values of $b=1+\delta$, $a=\frac{1}{2}$, $c=2+\delta$, where $\delta\coloneqq 10^{-5}$.
Simulation details
Our simulations utilize the state-of-the-art no-regret algorithm presented in Farina et al. (2022), the Log-Regularized Lifted Optimistic FTRL (hereafter, LRL-OFTRL). For each player, the algorithm uses the current gradient of the utility, the gradients from previous rounds, and a learning rate $\eta$ to decide which action to play in the next round. The important property of LRL-OFTRL is the logarithmic increase in regret as the dynamics progresses. Notably, the theoretical guarantee in Farina et al. (2022) applies only to concave games, and, to the best of our knowledge, there is currently no algorithm in the literature that guarantees sub-linear regret for non-concave games. [^9] In Appendix C.2 we show that indeed under non-concave ranking, the resulting dynamics when using LRL-OFTRL need not be no-regret dynamics. In our simulations, which are carried out in rounds, we define a publishers’ game $G$ and let the players choose their action in each round using the LRL-OFTRL algorithm. A simulation is said to converge if the average profile has reached an $\varepsilon$-NE. [^10] Interestingly, all of the simulations have converged not only in the average sense (which is guaranteed by our main result), but also in the stronger sense of last-iterate convergence. Throughout all simulations, we used homogeneous $\lambda_{i}=\lambda$ for all $i\in N$, the squared Euclidean semi-metric $d(x,y)=\frac{1}{k}||x-y||_{2}^{2}$, learning rate $\eta=1/2$ and $\varepsilon=10^{-4}$. In all simulations, we consider uniform demand distributions over a sample of $s\in\operatorname*{\mathbb{N}}$ information needs.111111 Note that while we restrict our simulations to uniform demand distributions, any demand distribution can be approximated with $s$ large enough.
@@
To evaluate the performance of the different ranking functions, we report three evaluation criteria: the publishers’ welfare, the users’ welfare and the convergence rate. [^12] In addition to these three evaluation criteria, in Appendix C.1 we analyze empirically the effect of the ranking function on the average regret and discuss an interesting correlation with the convergence rate. The two welfare measures are defined similarly to previous work (Madmon et al., 2023; Yao et al., 2024b). The publishers’ welfare is defined to be the sum of publisher utilities:
$\mathcal{U}(x)\coloneqq\sum_{i=1}^{n}u_{i}(x)=1-\lambda\sum_{i=1}^{n}d(x_{i},x_{0}^{i})$ (4)
@@
[…]
5.1 Comparison between ranking function families
Figure 1 presents several interesting trends. A prominent phenomenon in the figure is a trade-off between the different measures - whenever an activation function induces an equilibrium in which the publishers’ welfare is high, the users’ welfare is lower and the game dynamics converge more slowly. The activation functions that yield higher publishers’ welfare values are the ones that align the ranking function more closely with a uniform ranking (i.e., $r_{i}\equiv\frac{1}{n}$), or, in other words, the ones in which the impact of the publishers’ content on the ranking is less significant. When ranking the documents using these functions, the publishers are less incentivized to create relevant content and thus stay closer to their respective initial documents. This leads to higher publishers’ welfare and lower users’ welfare. On the contrary, there are less uniform ranking functions, in which the distances from the information needs have a greater influence on the ranking. Those ranking functions encourage publishers to compete for each information need and to generally make their documents more relevant to the users, which leads to higher users’ welfare and lower publishers’ welfare. [^14] To elucidate this point, let us compare the activation functions $g_{lin}$ and $g_{root}$. Note that $g_{root}$ is a concave transformation of the linear activation function. When applying a square root to a number in $[0,1]$, the smaller it is the bigger the factor it is increased by. Therefore, for any fixed profile, the ranking produced by $g_{root}$ is closer to uniform distribution than that produced by $g_{lin}$. In addition, when the ranking functions are less uniform, the absolute values of the gradients increase, which leads to faster convergence. In the rest of the section, we demonstrate that this order relation between the ranking functions is robust to changes in the ecosystem structure, such as changes in the game parameters, and show how these changes affect our evaluation criteria.
@@ 5.1 Comparison between ranking function families
The effect of the penalty factor $\lambda$
Figure 1 illustrates how $\lambda$, the penalty factor, influences our evaluation criteria, showing similar trends to those discussed in Madmon et al. (2023). One might expect an increase in $\lambda$ to harm the publishers’ welfare, since a larger $\lambda$ imposes greater penalty on the publishers for providing content that differs from their initial documents. Yet, a higher $\lambda$ also incentivizes the publishers to adhere more closely to their initial documents, which increases the publishers’ welfare. These two opposing forces explain the non-monotonic trend observed in Figure 1: for lower values of $\lambda$, the publishers’ welfare decreases as $\lambda$ increases, until it reaches a minimum point, after which it keeps increasing. The trend in the users’ welfare is simpler: it monotonically decreases as $\lambda$ increases, likely because higher penalties lead to less relevant content from publishers. Lastly, there is no significant impact of the penalty factor on the convergence rate.
@@
As depicted in Figure 3, increasing $s$ leads to an increase in the publishers’ welfare and a decrease in the users’ welfare, both trends eventually plateauing. A possible explanation is that when increasing $s$ while fixing $n$, a fixed number of publishers compete for an increasing number of information needs, which makes it harder to maximize exposure. Recall that a publisher’s utility from exposure is the expected value of her exposure for different information needs, so when a player is unable to meet a large portion of the information needs, she might be demotivated to try to be relevant to the users and focus on being faithful to her initial document, boosting the publishers’ welfare in the expense of the users. The convergence rate is affected by $s$ as well. When $s$ is small, the publishers must compete for the same information needs, which leads to more complicated dynamics and hence to slower convergence. As $s$ increases, the publishers tend to focus on different information needs (the ones closest to their initial document), therefore we get much simpler dynamics and faster convergence.
@@
5.2 Comparison within ranking function families
Beyond the comparison between the three ranking function families, tuning the hyperparameters within each family also reveals interesting phenomena. The results once again underscore the publisher-user trade-off discussed in §5.1: hyperparameter values that lead to higher publishers’ welfare result in lower users’ welfare and slower convergence rate, and vice versa. As was the case in the comparison between ranking function families, the hyperparameters’ values that lead to higher publishers’ welfare (and lower users’ welfare) are the values that make the ranking function closer to uniform ranking. For example, a decrease in the power parameter $a$ of $r^{root}$ can be seen in Figure 5 to correlate with an increase in publishers’ and a decrease in users’ welfare, and indeed when the power parameter $a$ approaches 00, the root PRF approaches a uniform ranking. That is, $r^{root}_{i}\to\frac{1}{n}$ as $a\to 0$. Similarly, Figures 6 and 7 demonstrate how both the intercept parameter $b$ of $r^{lin}$ and the shift parameter $c$ of $r^{log}$ are positively correlated with the publishers’ welfare and negatively correlated with the users’ welfare, as $r^{lin}_{i}\to\frac{1}{n}$ as $b\to\infty$ and $r^{log}_{i}\to\frac{1}{n}$ as $c\to\infty$.
Figure 5: The effect of $a$ in the root ranking function, with $\lambda=0.5$, $n=3,s=3$ and $k=3$.
Figure 6: The effect of $b$ in the linear ranking function, with $\lambda=0.5$, $n=3,s=3$ and $k=3$.
Figure 7: The effect of $c$ in the logarithmic ranking function, with $\lambda=0.5$, $n=3,s=3$ and $k=3$.
@@ 6 Discussion
We studied an information retrieval game in which publishers have initial documents they prefer to provide and initiated the analysis of no-regret dynamics in such games. We defined the class of PRFs, which are ranking functions that are determined by a one-variable scalar activation function, and established a full characterization of which PRFs induce socially-concave games and which induce concave games. We then conclude that any concave activation function guarantees the convergence of no-regret dynamics, thereby offering a system designer a rich family of ranking functions that, under the assumption that publishers minimize their regret, guarantee the reach of the system to a stable state. We then empirically investigated the publishers’ welfare and the users’ welfare, as well as the convergence rate, which measures how long it takes the ecosystem to reach stability. We examined how these criteria are affected by the ranking function and by changes in the ecosystem structure. In Appendix A we discuss several limitations of our work, which we believe can lead to significant future research directions in the field of learning dynamics of content creators within search and recommendation ecosystems.
Ethics and reproducibility statement
The authors of this paper declare that they do not find any potential violations of the ICLR code of ethics. All theoretical results of this paper (stated in §3 and §4) are presented with clearly stating any required assumptions. Rigorous proofs of all theoretical results can be found in Appendix B. All of the empirical results are presented with bootstrap confidence intervals (as mentioned in §5). The source code is available in the supplementary materials.
@@ Acknowledgments
This work was supported by funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement 740435). All authors contributed equally to this work. We would like to thank the anonymous reviewers for their helpful comments.
Appendix A Limitations
We now discuss several limitations of our work. Although we prove convergence of no-regret dynamics whenever the activation function is concave, our approach provides no insights regarding convergence in the non-concave case. One major challenge in studying no-regret dynamics under non-concave activation is the absence of an algorithm that guarantees sub-linear regret in such games (see Appendix C.2 for further discussion). We assumed each publisher produces a single document, while in practice, they may publish multiple documents to compete across various information needs. We also assumed a specific cost function to model the cost publishers incur when deviating from their initial documents. Another limitation of our experiments is that we drew $x_{0}$ and $x^{*}$ i.i.d from a uniform distribution, and the trends presented in our analysis are not guaranteed to hold for other ecosystem distributions. However, in Appendix C.3, we experiment with several distributions that relax these assumptions. A theoretical convergence rate analysis remains out of scope. Finally, we acknowledge the gap between our theoretical guarantee of convergence in the average sense and the empirical observations of last-iterate convergence. We believe that addressing these limitations in future work can provide valuable insights and enhance the applicability of our findings.
Appendix B Proofs
B.1 Proof of Lemma 1
@@
since $\sum_{i=1}^{n}r_{i}(x;x^{*})=1$ for any strategy profile $x\in X$ and information need $x^{*}\in[0,1]^{k}$. Now, $g(x)$ is concave since for every $i\in N$ and $x_{0}^{i}\in[0,1]^{k}$, $d(x_{i},x_{0}^{i})$ is convex in $x_{i}$. Regarding condition A2., it can be easily shown (using the linearity and monotonicity of expectation) that the convexity of $r_{i}(x;x^{*})$ in $x_{-i}$ for any $x_{i}\in X_{i},x^{*}\in[0,1]^{k}$ implies that $\mathbb{E}_{x^{*}\sim P^{*}}\left[r_{i}(x;x^{*})\right]$ is convex in $x_{-i}$ for any $x_{i}\in X_{i}$. Together with the fact that $d(x_{i},x_{0}^{i})$ is fixed in $x_{-i}$, we get that $u_{i}(x)=\mathbb{E}_{x^{*}\sim P^{*}}\left[r_{i}(x;x^{*})\right]-\lambda_{i}\cdot d(x_{i},x_{0}^{i})$ is convex in $x_{-i}$ for any $x_{i}\in X_{i}$. ∎
B.2 Proof of Theorem 2
Proof.
[…]
$r_{i}(x_{i},x_{-i};x^{*})=\frac{g\bigl{(}d(x_{i},x^{*})\bigr{)}}{\sum_{j=1}^{n}g\bigl{(}d(x_{j},x^{*})\bigr{)}}=\frac{C}{C+h(x_{-i})}$ (7)
$r_{i}(x_{i},x_{-i};x^{*})=\ell\bigl{(}h(x_{-i})\bigr{)}$ (8)
[…]
The above wrapped up the proof of the equivalence of the three conditions. We now proceed to prove the additional part of the Theorem, namely that if $d$ is strictly bi-convex, then $r$ induces strictly concave games if and only if $g$ is concave. Since we have already shown that if $r$ induces concave games then the activation is concave, it remains to show that if the activation is concave, then $r$ induces strictly concave games, which we now prove
Let $r$ be a proportional ranking function with concave activation $g$. Let $G$ be a game induced by $r$. We will show that $G$ is strictly concave by definition - fix some player $i\in N$ and let $x_{-i}\in X_{-i}$ be some strategy profile of all players but $i$. We need to show that $u_{i}(x_{i},x_{-i})$ is strictly concave in $x_{i}$. Let $x^{*}$ be an information need in the support of the demand distribution $P^{*}$ of $G$. Since $d$ is strictly bi-convex, $d(x_{i},x^{*})$ is strictly convex in $x_{i}$. Also, $g$ is concave, and recall that $g$ is assumed to be strictly decreasing (see Definition 3). Therefore, $g(d(x_{i},x^{*}))$ is strictly concave in $x_{i}$ as a composition of a concave and strictly decreasing function with a strictly convex function. Now, define the function:
$h(t)=\frac{t}{t+\sum_{j\neq i}g(d(x_{j},x^{*}))}$ (9)
is strictly concave in $x_{i}$ since the first term is strictly concave as an expectation of a strictly concave function and $\lambda_{i}\cdot d(x_{i},x_{0}^{i})$ is strictly convex in $x_{i}$ by the assumption that $d$ is strictly bi-convex. ∎
B.2.1 Full derivation of inequality $f^{\prime\prime}(\hat{a})>0$ from the proof of Theorem 2
This appendix provides a detailed derivation of the equation $f^{\prime\prime}(\hat{a})>0$ from the proof of the main result.
Proof.
By the definition of the utility in publishers’ games, and since in the game we constructed the distribution over the information needs is a one-hot distribution that places probability $1$ on $x^{*}=0$: